# LiRA: Linguistic Robust Anchoring for Cross-lingual Large Language Models

### v1

@@ 5.4 Ablation
The ablation results in Table 5 demonstrate the contribution of each component in LiRA. Removing the LLM Critic or Embeds Critic leads to the most significant performance drop, particularly on Pearson correlation and accuracy, highlighting the importance of dual-level critics for effective supervision. The translation and multilingual encoder modules also provide consistent gains, showing their role in enhancing cross-lingual generalization. Finally, eliminating the FIFO loss queue results in the largest degradation on nDCG@10, confirming its necessity in stabilizing optimization. Overall, each component is essential, and their synergy ensures the robustness of LiRA across tasks.
@@
Method nDCG@10 Pearson Accuracy
LiRA 77.71 75.00 89.2
   - w/o LLM Critic 71.29 72.19 87.7
   - w/o Embeds Critic 65.77 61.78 81.2
   - w/o Translations 75.48 74.39 86.4
   - w/o Multilingual encoder 75.59 72.43 85.3
   - w/o FIFO loss queue 64.29 69.82
@@ 6 Conclusion
We proposed LiRA, a framework for robust multilingual LLM adaptation that unifies retrieval, sentence ranking, and reasoning tasks under a common anchoring principle. By combining anchored representations with critic-guided alignment and queue-based objectives, LiRA consistently improves over strong Qwen3 baselines across both public benchmarks and our newly introduced LazRetrieval dataset. Ablation studies further validate the complementary contributions of each component. We hope our dataset and framework can inspire future work on multilingual LLM adaptation. In addition, our Ethics Statement and Reproducibility Statement are provided in the  A.5 and  A.6.
Appendix A Appendix
A.1 Theoretic Details
A.1.1 Math Proof
Theorem (Representation deviation bound).
Under Assumptions 12 and Definitions 12, let the optimal English representation be ${\bm{z}}^{\star}=[\,h(y^{\star});\;h(y^{\star})\,]$ and the framework output be ${\bm{z}}=[\,g(x);\;h(y)\,]$. Then
$\bigl\|{\bm{z}}-{\bm{z}}^{\star}\bigr\|_{2}\;\leq\;\epsilon_{1}\;+\;C\,\sqrt{2\,\epsilon_{2}}\,$, (13)
@@ A.1.1 Math Proof
Let $y^{\star}$ be a perfect translation such that $p(s\mid x)=p(s\mid y^{\star})$. By block structure and the triangle inequality,
@@
Using the RKHS mean-embedding view of $h$ (Definition 1) and the bounded-kernel assumption (see, e.g., Sriperumbudur et al. (2010)),
@@
Plugging equation 17 into equation 15 and then into equation 14 yields
@@
$\bigl\|f_{\mathrm{LLM}}({\bm{z}})-f_{\mathrm{LLM}}({\bm{z}}^{\star})\bigr\|_{2}\;\leq\;L^{\mathrm{loc}}(y;\delta)\,\Bigl(\,\epsilon_{1}+C\,\sqrt{2\,\epsilon_{2}}\,\Bigr).$ (18)
@@
By the (local) Lipschitz property of $f_{\mathrm{LLM}}$ and Theorem Theorem,
$\bigl\|f_{\mathrm{LLM}}({\bm{z}})-f_{\mathrm{LLM}}({\bm{z}}^{\star})\bigr\|_{2}\;\leq\;L^{\mathrm{loc}}(y;\delta)\,\|{\bm{z}}-{\bm{z}}^{\star}\|_{2}\;\leq\;L^{\mathrm{loc}}(y;\delta)\,\Bigl(\epsilon_{1}+C\sqrt{2\,\epsilon_{2}}\Bigr).$
@@
$\bigl\|f_{\mathrm{LLM}}({\bm{z}})-f_{\mathrm{LLM}}({\bm{z}}^{\star})\bigr\|_{1}\;\leq\;0.034\cdot\Bigl(\,\epsilon_{1}+1.9423\,\sqrt{\epsilon_{2}}\Bigr).$ (19)
A.1.2 About Definition
@@
Remark (On estimating the boundedness constant $C$).
@@
The data-local Lipschitz constant
$L_{h}^{\mathrm{loc}}(y;\delta)=\max_{y^{\prime}\in\mathcal{N}_{\delta}(y)}\frac{\bigl\|h(y)-h(y^{\prime})\bigr\|_{2}}{d_{\mathrm{tok}}(y,y^{\prime})}$ (25)
@@
  • $h:\mathcal{Y}\!\to\!\mathbb{R}^{d}$ is any fixed sentence encoder (e.g., Qwen-3, BERT).
@@
A.2 Dataset
@@ A.3 Experimental Details
Arca Translation Selections.
As shown in Figure 10. On mlqaRetrieval, mgsm, and belebeleRetrieval, Qwen3-32B is the most frequently selected candidate (29.4%, 34.0%, and 33.8%, respectively); together with DeepSeek-R1-Distill-Qwen-32B (27.0%, 31.8%, 21.4%), the Qwen-family accounts for the majority of selections (56.4%, 65.8%, and 55.2%). On LazRetrieval, the distribution is more balanced with Llama-33-70B-Instruct at 27.2%, gemma-2-27b-it at 26.8%, and the Qwen-family totaling 46.0%. Across datasets, Arca shows a consistent preference toward Qwen-family candidates (Qwen3 or the Qwen-distilled DeepSeek variant), especially on mgsm where they comprise 65.8% of selections. A plausible explanation is architectural alignment and feature-space compatibility with our frozen evaluator, which is Qwen3-72B. Using the same family for evaluation can introduce a mild inductive bias that favors stylistic and semantic choices characteristic of that architecture. We therefore report these breakdowns to make the potential bias explicit and to encourage future work to cross-check with evaluators from different families.
@@
STS22 mT5-XL Qwen3-Embedding-8B 5k 1 10 4
BelebeleRetrieval mT5-XL Qwen3-Embedding-8B 1k 1 5 8
@@
MGSM mT5-XL Qwen3-14B 4 2 - 1
X-CSQA mT5-XL Qwen3-14B 4 2 - 1
@@
A.4 Model Details
Backbones.
Multilingual encoder: mT5-XL (encoder only; decoder frozen). LLM Evaluator (frozen): Qwen/Qwen3-70B is used only for evaluation/critique; all its parameters are kept frozen. LaSR (trainable): the LaSR module is trained end-to-end during our experiments (gradients do not propagate into the frozen evaluator). When needed, token representations are read via get_input_embeddings() (no gradient flow).

### v2

@@ 5.4 Ablation
The ablation results in Table 5 demonstrate the contribution of each component in LiRA. Removing the LLM Critic or Embeds Critic leads to the most significant performance drop, particularly on Pearson correlation and accuracy, highlighting the importance of dual-level critics for effective supervision. The translation and multilingual encoder modules also provide consistent gains, showing their role in enhancing cross-lingual generalization. Finally, eliminating the FIFO loss queue results in the largest degradation on nDCG@10, confirming its necessity in stabilizing optimization. Overall, each component is essential, and their synergy ensures the robustness of LiRA across tasks. In our ablation study, the three reported metrics are obtained on different tasks: nDCG@10 is evaluated on LazRetrieval, Pearson is evaluated on STS22, and Acc. is evaluated on MGSM.
@@
Method nDCG@10 Pearson Acc.
LiRA (full) 77.71 75.00 71.1
  $\hookrightarrow$ – LLM critic 71.29 72.19 68.9
  $\hookrightarrow$ – Embeds critic 65.77 61.78 67.3
  $\hookrightarrow$ – Translations 75.48 74.39 70.5
  $\hookrightarrow$ – Multi-enc 75.59 72.43 69.5
  $\hookrightarrow$ – FIFO queue 64.29 69.82
5.5 Supplementary Experiments
To offer additional insight into the practical behavior of our framework, we provide supplementary analyses in the appendix. In particular, we report (i) an empirical breakdown of Arca’s translation candidate selections across datasets (Appendix C, Figure 10), which helps reveal potential evaluator-induced preferences; (ii) estimates of the RKHS boundedness surrogate $\widehat{C}$ across backbone scales (Appendix A, Table 6); and (iii) data-local Lipschitz statistics $L_{\mathrm{emp}}(\delta)$ under token-level perturbations (Appendix A, Table 7). Together, these results provide practical guidance for instantiating our theoretical bounds and interpreting model behavior beyond the primary benchmarks.
@@ 6 Conclusion
We proposed LiRA, a framework for robust multilingual LLM adaptation that unifies retrieval, sentence ranking, and reasoning tasks under a common anchoring principle. By combining anchored representations with critic-guided alignment and queue-based objectives, LiRA consistently improves over strong Qwen3 baselines across both public benchmarks and our newly introduced LazRetrieval dataset. Ablation studies further validate the complementary contributions of each component. We hope our dataset and framework can inspire future work on multilingual LLM adaptation.
Appendix A Theoretic Details
A.1 Math Proof
Theorem 1 (Representation deviation bound).
Under Assumptions 12 and Definitions 12, let the optimal English representation be $\mathbf{z}^{\star}=[\,h(y^{\star});\;h(y^{\star})\,]$ and the framework output be $\mathbf{z}=[\,g(x);\;h(y)\,]$. Then
$\bigl\|\mathbf{z}-\mathbf{z}^{\star}\bigr\|_{2}\;\leq\;\epsilon_{1}\;+\;C\,\sqrt{2\,\epsilon_{2}}\,$, (10)
@@ A.1 Math Proof
Let $y^{\star}$ be a ideal translation such that $p(s\mid x)=p(s\mid y^{\star})$. By block structure and the triangle inequality,
@@
Using the RKHS mean-embedding view of $h$ (Definition 1) and the bounded-kernel assumption (see, e.g., (Sriperumbudur et al., 2010)),
@@
Plugging (14) into (12) and then into (11) yields
@@
By the (local) Lipschitz property of $f_{\mathrm{LLM}}$ and Theorem 3.2,
$\bigl\|f_{\mathrm{LLM}}(\mathbf{z})-f_{\mathrm{LLM}}(\mathbf{z}^{\star})\bigr\|_{2}\;\leq\;L^{\mathrm{loc}}(y;\delta)\,\|\mathbf{z}-\mathbf{z}^{\star}\|_{2}\;\leq\;L^{\mathrm{loc}}(y;\delta)\,\Bigl(\epsilon_{1}+C\sqrt{2\,\epsilon_{2}}\Bigr).$
@@
$\bigl\|f_{\mathrm{LLM}}(\mathbf{z})-f_{\mathrm{LLM}}(\mathbf{z}^{\star})\bigr\|_{1}\;\leq\;0.034\cdot\Bigl(\,\epsilon_{1}+1.9423\,\sqrt{\epsilon_{2}}\Bigr).$ (16)
A.2 About Definition
@@
Remark 1 (On estimating the boundedness constant $C$).
@@
The data-local Lipschitz constant is defined w.r.t. the downstream scorer $f_{\mathrm{LLM}}$ as
@@
  • $f_{LLM}$: is any fixed downstream model (e.g., Qwen-3, Qwen-3-Embedding, BERT).
@@
A.3 Why Concatenate TWO Representation Paths?
$g(x)=s+\eta_{g},\qquad h(y)=s+\eta_{h},$
Define the information gain of concatenation:
$\Delta I\;=\;I\!\bigl(s;\,[g(x),h(y)]\bigr)\;-\;I\!\bigl(s;\,g(x)\bigr),$
$I\!\bigl(s;\,[g,h]\bigr)$ $=I(s;g)+I(s;h\mid g)$
$=I(s;g)+H(s\mid g)-H(s\mid g,h)$
$\geq I(s;g),$ (24)
with equality iff $s\!\perp\!\!\!\perp\!h\mid g$ (i.e., $h(y)$ provides no additional semantic information beyond what is already contained in $g(x)$). In cross-lingual settings, translation noise $\eta_{h}$ and anchoring noise $\eta_{g}$ are complementary, hence $I(s;h\mid g)\!>\!0$ and $\Delta I\!>\!0$. Therfore, if $\mathrm{Var}(\eta_{g,k})\!\to\!\infty$ on some dimension $k$ (e.g., severe LRL ambiguity), then
$I(s;g)\;\leq\;\tfrac{1}{2}\log\!\Bigl(1+\tfrac{\sigma_{s,k}^{2}}{\mathrm{Var}(\eta_{g,k})}\Bigr)\!\to 0.$
Appendix B Dataset
@@ Appendix C Experimental Details
C.1 Arca Translation Selections-Equipped LiRA-Max
As shown in Figure 10. On mlqaRetrieval, mgsm, and belebeleRetrieval, Qwen3-32B is the most frequently selected candidate (29.4%, 34.0%, and 33.8%, respectively); together with DeepSeek-R1-Distill-Qwen-32B (27.0%, 31.8%, 21.4%), the Qwen-family accounts for the majority of selections (56.4%, 65.8%, and 55.2%). On LazRetrieval, the distribution is more balanced with Llama-33-70B-Instruct at 27.2%, gemma-2-27b-it at 26.8%, and the Qwen-family totaling 46.0%. Across datasets, Arca shows a consistent preference toward Qwen-family candidates (Qwen3 or the Qwen-distilled DeepSeek variant), especially on mgsm where they comprise 65.8% of selections. A plausible explanation is architectural alignment and feature-space compatibility with our frozen evaluator, which is Qwen3-32B. Using the same family for evaluation can introduce a mild inductive bias that favors stylistic and semantic choices characteristic of that architecture. We therefore report these breakdowns to make the potential bias explicit and to encourage future work to cross-check with evaluators from different families.
@@ C.1 Arca Translation Selections-Equipped LiRA-Max
STS22 mT5-XL Qwen3-Embedding-8B 8 1 10 4
BelebeleRetrieval mT5-XL Qwen3-Embedding-8B 8 1 5 8
@@
MGSM mT5-XL Qwen3-8B 4 2 - 1
X-CSQA mT5-XL Qwen3-8B 4 2 - 1
@@
C.2 Prompt Engineer
Our translation-aware training relies on a translator module and a critic module. In practice, however, off-the-shelf LLM translators may produce non-translation artifacts (e.g., meta prefixes such as “Here is the translation:”, unnecessary politeness, or extra explanations), which introduce avoidable noise into both the translation candidates and the downstream scoring signals. To reduce such artifacts and to make the translation outputs more consistent across languages, we adopt a simple but effective prompt-engineering strategy for both translation and evaluation.
Translation prompt. We instruct the model to act as a professional translator and output only the translated text without any additional commentary:
You are a professional translator, proficient in various languages. Please translate the following phrase or sentence into English: ’{text}’. Output only the translation results without any other explanations.
Critic prompt. To compare multiple translation candidates, the critic evaluates each candidate on three dimensions—semantic fidelity, emotional consistency, and pragmatic tone—and is required to output strict JSON to avoid verbose judgments that are hard to parse:
Please strictly evaluate the following translation on three dimensions:
  • Semantic Fidelity (1–10): Accuracy in conveying original meaning
  • Emotional Consistency (1–10): Alignment with original emotional tone
  • Pragmatic Tone (1–10): Appropriateness in contextual usage and style
You MUST only output JSON format with keys ’semantic’, ’emotional’, ’pragmatic’ and values 1–10, without any other explanations.
Origin text in {lang}: ’{sentence}’
Translation in En: ’{translation}’
This design serves two purposes: (i) it suppresses extraneous natural-language responses that otherwise contaminate translation candidates and destabilize training, and (ii) it yields structured, dimension-wise scores that can be directly integrated into the actor–critic learning signal without additional heuristics.
C.3 Bad Case Analysis
Despite strong translation ability, large language models frequently inject small but non-negligible noise into translation outputs, especially when the input is long, contains news-style formatting, or includes imperative phrases and boilerplate (e.g., “Follow us on Telegram”). A common failure mode is meta-text leakage: the model prepends or appends non-translation content such as “Sure, here is the translation:”, “Here is the translated sentence:”, or polite closing statements like “If you need any other help, I’m glad to help you.” Although such artifacts are harmless for human readers, they are problematic in our setting because they (i) change the token distribution of the “translation” string, (ii) inject irrelevant stylistic cues that the encoder may latch onto, and (iii) distort critic scoring by mixing translation quality with adherence-to-instruction behavior. We provide a representative example below. Given a long-form input, some LLM translators return a translation wrapped with assistant-style meta prefixes and extra explanations, instead of outputting the translation alone:
Expected translation candidate:
Baku, Azerbaijan, Jan. 1 ... ‘‘2019 will go down in history as a year of in-depth reforms ...’’
Noisy translation candidate:
Sure, here is the translated sentence:
‘‘Baku, Azerbaijan, January 1 ... ‘‘2019 will go down in history as a year of significant reforms ...’’ ...’’
Such meta-text leakage constitutes translation noise under our theoretical framing (semantic drift and formatting artifacts), and it can propagate into both representation learning and the actor–critic reward signal. Concretely, the added preambles (e.g., “Sure, here is …”) introduce irrelevant tokens and style markers that are absent from genuine translations, making the translation string less comparable across candidates and languages.
Our prompt-engineering strategy mitigates this issue by enforcing output-only translation for the translator and JSON-only scoring for the critic. Empirically, this reduces the rate of meta-text leakage and yields more uniform translation candidates in style and formatting, leading to a cleaner training signal and more stable translation-aware optimization, especially for long-form inputs with news-style templates.
@@ C.3 Bad Case Analysis
Appendix D Model Details
D.1 About Model
Multilingual encoder: mT5-XL (encoder only; decoder frozen). LLM Evaluator (frozen): Qwen/Qwen3-32B is used only for evaluation/critique; all its parameters are kept frozen. LaSR (trainable): the LaSR module is trained end-to-end during our experiments (gradients do not propagate into the frozen evaluator). When needed, token representations are read via get_input_embeddings() (no gradient flow).
@@ D.1 About Model
D.2 Measuring $\epsilon_{1}$ and $\epsilon_{2}$
Our theory only assumes: (i) there exists a mapping error between representation vectors, and (ii) machine translation may introduce noise. Both phenomena are widely observed in cross-lingual alignment and MT-based transfer settings. We introduce an “ideal” representation vector $\mathbf{z}^{\star}$ that corresponds to the noisy, error-prone representation $\mathbf{z}$ purely as a mathematical construct, rather than as a strict requirement on real-world conditions. Importantly, the purpose of these assumptions is to derive and justify our training objective: Assumptions 1–2 are directly aligned with our loss, and $\mathbf{z}^{\star}$ serves to emphasize that errors arise along two orthogonal dimensions—semantic drift (translation distortion) and vector mapping mismatch (representation deviation).
To make these assumptions empirically verifiable, we estimate both error terms using observable proxies during training. For the representation mapping error $\epsilon_{1}$, we track the mismatch between the feature-path representation and the translation-path representation:
$\hat{\epsilon}_{1}~:=~\mathbb{E}_{x}\Big[\big\|g(x)-h(\hat{y}(x))\big\|_{2}\Big],$ (25)
where $g(x)$ denotes the feature-path embedding and $h(\hat{y}(x))$ denotes the translation-path embedding produced by the selected translation $\hat{y}(x)$ (e.g., the actor/critic-chosen candidate). In addition, we report the cosine similarity $\cos(\mathbf{z},\mathbf{z}^{\star})$ as a normalized proxy to visualize the contraction of representation deviation across epochs (higher is better).
For the translation distortion $\epsilon_{2}$, since the ideal translation $y^{\star}$ is not observable, we use a semantic-divergence proxy that captures the degree of drift introduced by translation:
$\hat{\epsilon}_{2}~:=~\mathbb{E}_{x}\Big[1-\cos\big(\mathrm{Emb}(x),\mathrm{Emb}(\hat{y}(x))\big)\Big],$ (26)
In Eq. (26), we instantiate $\mathrm{Emb}(\cdot)$ using the frozen encoder states of $f_{\text{llm}}$ (i.e., the LLM representations before decoding). Concretely, given an input text $u$ (either the source sentence $x$ or the selected translation $\hat{y}(x)$), we obtain token-level hidden states $\mathbf{H}_{u}=\mathrm{Enc}_{f_{\text{llm}}}(u)\in\mathbb{R}^{T\times D}$, and compute a sentence-level embedding by masked mean pooling:
$E_{f_{\text{llm}}}(u)~=~\frac{1}{\sum_{t}m_{t}}\sum_{t=1}^{T}m_{t}\,\mathbf{H}_{u}[t],$ (27)
where $m_{t}\in\{0,1\}$ is the attention mask. We keep $f_{\text{llm}}$ frozen and use a fixed prompting/formatting template for all inputs in this appendix to ensure $E_{f_{\text{llm}}}(\cdot)$ is a stable diagnostic signal.
Accordingly, our translation-distortion proxy is measured as
$\hat{\epsilon}_{2}~:=~\mathbb{E}_{x}\Big[1-\cos\big(E_{f_{\text{llm}}}(x),\,E_{f_{\text{llm}}}(\hat{y}(x))\big)\Big].$ (28)
Lower $\hat{\epsilon}_{2}$ indicates better semantic faithfulness and reduced translation noise (as measured in the representation space of $f_{\text{llm}}$).
Tab. 11 reports the evolution of the similarity between $\mathbf{z}$ and $\mathbf{z}^{\star}$ (cosine similarity proxy) and the corresponding training loss across epochs on three representative Arca training tasks. We observe a consistent increase in similarity and a monotonic decrease in loss, which is consistent with the theoretical interpretation that the representation deviation contracts as training proceeds. Tab. 12 summarizes the measured diagnostics $\hat{\epsilon}_{1}$ and $\hat{\epsilon}_{2}$ under different training settings.
Table 11: Similarity between $\mathbf{z}$ and $\mathbf{z}^{\star}$ (cosine similarity proxy) and training loss across epochs (averaged over three Arca training tasks). Higher similarity and lower loss indicate improved anchoring and reduced deviation.
Epoch 1 2 3 4 5 6 7 8 9 10
Sim. 0.01 0.14 0.34 0.45 0.54 0.60 0.64 0.67 0.69 0.71
Loss 1.94 1.23 0.97 0.74 0.62 0.54 0.42 0.35 0.29 0.21
Table 12: Diagnostics for representation mapping error and translation distortion. Lower is better for $\hat{\epsilon}_{1}$ and $\hat{\epsilon}_{2}$. Fill in with your measured values (mean $\pm$ std if available).
Setting Task $\hat{\epsilon}_{1}\downarrow$ $\hat{\epsilon}_{2}\downarrow$ Sim. $\uparrow$ nDCG@10 $\uparrow$
Qwen3-Embedding-8B LazRetrieval 0.81 0.54 0.01 69.4
+Arca LazRetrieval 0.19 0.23 0.78 69.9
+Arca + LaSR LazRetrieval 0.19 0.23 0.78 71.1
D.3 A more complex training pipeline?
While LiRA introduces translation-aware training, our pipeline is designed to be budget-flexible and engineering-friendly. First, translations can be prepared offline in advance, which substantially reduces online training overhead. Second, the translator can be replaced by a smaller MT model when compute is constrained. We report resource usage for both training and inference in Tab. 13 and Tab. 14, measured in single A100-80G GPU hours. [^1] The reported “GPU hours” measure the end-to-end wall-clock GPU time under our implementation setup.
Here, pass@k indicates that $k$ LLM translators are used along the forward path (i.e., $k$ translation candidates are produced by LLMs); pass@0 uses only MT models. In practical deployment, one may combine a lightweight MT model with an LLM (or use MT only). Notably, even pass@0 already outperforms the original Qwen3-Embedding-8B baseline in our experiments, indicating that LiRA provides a principled way to customize computation budgets and select translator capacity according to available resources.
Table 13: Training resource usage (single A100-80G GPU hours). pass@k denotes using $k$ LLM translators along the forward path.
Setting Pass@4 Arca Pass@4 LaSR Pass@2 Arca Pass@2 LaSR Pass@0 Arca Pass@0 LaSR
Offline translation 0.15h 0.20h 0.15h 0.20h 0.15h 0.20h
Online translation 1.50h 2.00h 0.80h 1.10h 0.30h 0.40h
Table 14: Inference resource usage (single A100-80G GPU hours) under different pass@k settings.
Setting pass@4 pass@2 pass@0
Offline translation 0.30h 0.30h 0.30h
Online translation 3.80h 2.10h 0.90h
Pass@k.
We evaluate with a $k$-way budget: pass@k is the number of LLM translators used in one LiRA forward pass (k=0k{=}0 uses only MT; k>0k{>}0 follows Table 15). Table 16 aggregates MLQA Retrieval, BelebeleRetrieval, and STS22. Across $k\in\{0,1,2,3,4\}$, LiRA outperforms the baseline on all three datasets. Both models improve as $k$ increases, with diminishing gains beyond k=2k{=}2 and a small dip at k=3k{=}3 on MLQA. Even at k=0k{=}0, LiRA beats the baseline, showing a tunable budget–quality trade-off without large translators at inference.
Table 15: Translator configurations for pass@k (abbreviations: OPUS = OPUS-MT; M2M = m2m100; N600 = nllb-200-600M; N3B = nllb-200-3.3B; DS-R1 = Deepseek-R1-Distill-Qwen-32B).
pass@k T1 T2 T3 T4
pass@0 OPUS M2M N600 N3B
pass@1 Llama OPUS M2M N3B
pass@2 Llama Gemma M2M N3B
pass@3 Llama Gemma Qwen N3B
pass@4 Llama Gemma Qwen DS-R1
Table 16: Combined pass@k results on three datasets (higher is better).
Method pass@k MLQA-R Bele-R STS22
Qwen3-8B pass@0 79.96 82.27 69.64
pass@1 80.41 83.54 70.01
pass@2 80.45 83.97 70.72
pass@3 80.79 84.55 71.32
pass@4 81.13 85.94 71.64
With LiRA pass@0 81.15 86.00 73.01
pass@1 81.56 86.15 73.54
pass@2 81.79 86.67 74.11
pass@3 81.53 86.69 74.39
pass@4 82.01 87.03 75.00
D.4 Objective